OpenStudy (anonymous):

find the height of a prism with a base in the form of the sector in which the central anlge is 25 degrees. its volume is found to be 350 inch, cube and the height is four times the radius of the base

OpenStudy (anonymous):

volume of prism = (base area) (height)

OpenStudy (anonymous):

how sir ???

OpenStudy (anonymous):

base area is fraction of circle = (25/360) pi r^2 h = 4 r solve volume = (base area) (height)

OpenStudy (anonymous):

and then sir ???

OpenStudy (anonymous):

350 = (25/360) pi r^2 (4 r) solve for r

OpenStudy (anonymous):

r^3 = etc. r = cube root (etc.)

OpenStudy (anonymous):

i will get r ??